# Learning Neural Free-Energy Functionals with Pair-Correlation Matching

### v3

@@ Abstract
The intrinsic Helmholtz free-energy functional, the centerpiece of classical density functional theory (cDFT), is at best only known approximately for 3D systems. Here we introduce a method for learning a quasi-exact neural-network approximation of this functional by exclusively training on a dataset of radial distribution functions, circumventing the need to sample costly heterogeneous density profiles in a wide variety of external potentials. For a supercritical 3D Lennard-Jones system, we demonstrate that the learned neural free-energy functional accurately predicts planar inhomogeneous density profiles under various complex external potentials obtained from simulations.
Figure 1: 1. Bulk densities in planar geometry $\rho(z_{i})=\rho_{b}$ and radial distribution functions $g(r)$ are sampled from Monte Carlo simulations of homogeneous bulk systems of Lennard-Jones particles. Each $g(r)$ is converted to the second functional derivative of the excess free-energy $\delta^{2}\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})$ by employing the Ornstein-Zernike equation. 2. Through automatic differentiation (autodiff2), the neural functional $F_{\theta}^{(2)}$ is optimized to fit the Hessian of the model output with respect to input density profiles to $\delta^{2}\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})$. 3. The optimized model can then be applied in cDFT to obtain non-uniform equilibrium density profiles through automatic differentiation (autodiff) and the free energy $F_{\theta}^{(2)}$ for a system of Lennard-Jones particles subjected to arbitrary external potentials.
@@
In this Letter, we introduce a neural free-energy functional that we train using pair-correlation matching, a novel optimization objective that matches the Hessian of the neural approximation to pair correlations of particles. We show that pair-correlation matching yields a neural functional that accurately predicts the complete excess free energy for a one-component system of interacting particles. The differentiable nature of this neural functional allows access to, and learning from, various structural and thermodynamic properties by utilizing the first and second functional derivatives of $\mathscr{F}_{\text{exc}}[\rho]$. Unlike previous ML approaches to cDFT [11, 12, 13, 14], our neural functional is trained by directly learning pair correlations from radial distribution functions sampled from short simulations of homogeneous bulk systems (illustrated in Fig. 1), rather than inferring from a costly dataset of inhomogeneous densities. We demonstrate that this neural free-energy functional can be applied in the cDFT framework to achieve accurate estimates for inhomogeneous density profiles in external fields without ever having seen any inhomogeneous densities during training. Simultaneously, the neural free-energy functional provides accurate estimates of the excess free energy.
@@
with $\beta=1/k_{B}T$, $\Lambda$ the thermal wavelength, $\mu$ the chemical potential and $V_{\text{ext}}(\mathbf{r})$ the external potential. This self-consistency relation can be leveraged to find $\rho_{0}(\mathbf{r})$ through recursive iteration. Since $\rho_{0}(\mathbf{r})$ is dependent on $\mathscr{F}_{\text{exc}}[\rho]$ through the first functional derivative $\delta\mathscr{F}_{\text{exc}}[\rho]/\delta\rho(\mathbf{r})$, previous approaches to leveraging machine learning for cDFT [11, 12, 13] involve training a model to capture $\delta\mathscr{F}_{\text{exc}}/\delta\rho(\mathbf{r})$, which can be derived from sampled inhomogeneous equilibrium density profiles and employing Eq.(1).
@@
Figure 2: Evaluation of neural free-energy functionals $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$, where $F_{\theta}^{(1)}$ is optimized by matching inhomogeneous one-body densities and $F_{\theta}^{(2)}$ by pair-correlation matching in the homogeneous bulk. a) Density profiles of a Lennard-Jones system in a planar geometry characterized by an external potential (shown in gray) at a chemical potential of $\beta\mu=0$ obtained from DFT using $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ and the mean-field approximation $F_{\text{exc}}^{\text{MF}}$, along with the simulated density profile. b) Comparison of the free-energy estimates using $F_{\theta}^{(1)}$, $F_{\theta}^{(2)}$, and $F_{\text{exc}}^{\text{MF}}$ for the specific external potential shown in a), for the same chemical potential of $\beta\mu=0$. c) Integrated absolute error (IAE) between density profiles obtained from DFT using $F_{\theta}^{(1)}$, $F_{\theta}^{(2)}$, and $F_{\text{exc}}^{\text{MF}}$, along with the densities sampled from simulation. d) Absolute error of the excess free energy from DFT and simulations. c)/d) Data is shown for 150 distinct external potentials, evaluated across the range $-4<\beta\mu<6$, with steps of $\Delta\beta\mu=0.1$. The area of the mean $\pm$ standard deviation is colored. The error is shown up to the point where the DFT iterations stop to converge to a solution within 1000 iterations. e) The pressure and the chemical potential obtained from DFT and simulations. f) The laterally integrated direct correlation function $\bar{c}_{b}^{(2)}(z)$ at $\rho_{b}\sigma^{3}=0.62$ and the radial distribution function $g(r)$ obtained from simulation and DFT.
@@
In this work, we train a convolutional neural network to directly learn the excess free energy $\mathscr{F}_{\text{exc}}[\rho]$, enabling the straightforward calculation of functional derivatives by (auto-)differentiating the neural functional with respect to its inputs. We focus on 3D systems in a planar geometry, where the excess free energy of a system of area $A$ is a functional of the density $\rho(z)$, which is constant across any plane parallel to the $xy$-plane, i.e., $\rho(z)=\rho(x,y,z)$ with $\rho(x,y,z)=\rho(x^{\prime},y^{\prime},z)$ for all $(x,y),(x^{\prime},y^{\prime})$ within the confines of $A$. We represent the excess free-energy functional $\mathscr{F}_{\text{exc}}[\rho]$ as a neural network $F_{\theta}^{(2)}(\rho_{1},\cdots\rho_{n})$, with $\rho_{i}$ the density at grid point $z_{i}$ for $i\in\{1,\ldots,n\}$, network parameters $\theta$, and where the upper index ‘$(2)$’ indicates that the neural network is optimized for pair correlations of bulk systems obtained from simulations.
@@
with $\hat{h}(k)$ the Fourier transform of the total correlation function $h(r)=g(r)-1$ [8, 17]. Thus we train our neural network such that its Hessian $\partial^{2}F_{\theta}^{(2)}/\partial\rho_{i}\partial\rho_{j}$ represents well the laterally integrated direct correlation function obtained from homogeneous bulk simulations [17]. We will refer to this approach as pair-correlation matching.
@@
We illustrate the power of this methodology for systems interacting with a Lennard-Jones potential truncated at $r_{\text{cut}}=4\sigma$ with $\sigma$ the particle diameter, shifted upwards by $\epsilon_{\text{cut}}=0.98\cdot 10^{-3}\epsilon$ with $\epsilon$ the well depth, at a temperature $k_{B}T/\epsilon=2$, i.e. above the critical point. Setting $\Lambda=\sigma$ throughout, we perform grand-canonical Monte Carlo (GCMC) simulations of homogeneous bulk systems at a variety of chemical potentials $\beta\mu\in[-4,0.5]$, resulting in bulk densities $\rho_{b}\sigma^{3}\in[0.02,0.67]$. We employ $10^{9}$ trial moves in a cubic box with an edge length $L=10\sigma$ (hence with area $A=L^{2}$ and volume $V=L^{3}$) and apply periodic boundary conditions. We sample $g(r)$ and $\rho_{b}$ and convert each $g(r)$ to $\bar{c}_{b}^{(2)}(z)$ using Eqs. (2) and (3) with a grid-spacing $\Delta z=\sigma/32$. We create a data set of 800 combinations of $\bar{c}_{b}^{(2)}(z)$ and $\rho_{b}$, from which we learn the parameters $\theta$ of $F_{\theta}^{(2)}$. The neural network architecture of $F_{\theta}^{(2)}$ is a six-layer convolutional neural network with periodic padding [18] to match the system’s periodic boundary conditions. Each convolution uses a kernel size of 3 with a dilation rate of 3. The number of channels per hidden layer is configured as follows: $[16,16,32,32,64,64]$, such that $\theta$ consists of 24.4K parameters [17].
@@
To evaluate the accuracy of our neural excess free-energy functional $F_{\theta}^{(2)}$, we compare it to the Van der Waals-like mean-field approximation $F_{\text{exc}}^{\text{MF}}$, which treats the attractions of Lennard-Jones particles as a perturbation on the hard-sphere system, as implemented in PyDFTlj [1]. We use the White-Bear mark II version of FMT for the excess free energy of the hard-sphere system [19]. Additionally, we compare to $F_{\theta}^{(1)}$, which is a neural functional trained by minimization of the error between $(1/\Delta z)\partial F^{(1)}_{\theta}/\partial\rho_{i}$ and $\delta\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})$ rather than by pair-correlation matching. This neural functional is trained on a dataset of 800 non-uniform densities, subjected to the same set of chemical potentials as before [17]. By approximating $\delta\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})$ by the gradient $(1/\Delta z)\partial F^{(n)}_{\theta}/\partial\rho_{i}$ for $n=1$ and 2, both neural functionals are applied in Picard iterations [20, 21, 22] to obtain DFT estimates for the equilibrium density profiles of inhomogeneous systems according to Eq. (1).
@@
The DFT results for an exemplary external potential at $\beta\mu=0$ are shown in Fig. 2a, where we observe that the neural functionals $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ provide similar estimates, both outperforming $F_{\text{exc}}^{\text{MF}}$. For the same external potential, we evaluate the accuracy of DFT estimates for the free energy for a range of chemical potentials $-4<\beta\mu<6$ (Fig. 2b). We compare with the excess free energy obtained from GCMC simulations through thermodynamic integration [17]. We observe that both neural functionals outperform $F_{\text{exc}}^{\text{MF}}$ within the range of $\mu$ values in the training set, exhibiting good agreement with the simulations. The DFT estimates are shown until the DFT iterations diverge and errors become unmanageable. This reveals that $F_{\theta}^{(1)}$ diverges rapidly when extrapolating beyond the training set, even earlier than $F^{\text{exc}}_{\text{MF}}$. In contrast, $F_{\theta}^{(2)}$ is capable of converging to a solution far beyond the trained $\mu$ range.
For a more detailed comparison of the performance of the free-energy functionals for various inhomogeneous systems, we performed separate DFT calculations for 150 distinct external potentials, evaluated across the range $-4<\beta\mu<6$. For both the density estimates (Fig. 2c) and the free-energy estimates (Fig. 2d), we observe excellent agreement between the $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ functionals and simulated data. They demonstrate similar performance for $\mu$ values within the training set range, and outperform the $F_{\text{exc}}^{\text{MF}}$ functional for all evaluated external potentials (Fig. 2). Again, we observe that $F_{\theta}^{(2)}$ surpasses $F_{\theta}^{(1)}$ when extrapolating beyond the $\mu$ range encountered during training.
@@
and the chemical potential $\mu$ following Eq. (1). Again, we find excellent agreement of both neural functionals with simulations within the training set, and superior agreement of $F_{\theta}^{(2)}$ at higher densities, as shown in Fig. 2e. Lastly, we demonstrate that $F_{\theta}^{(2)}$ compared to both $F_{\theta}^{(1)}$ and $F^{\text{exc}}_{\text{MF}}$ provides accurate estimates for the laterally integrated direct correlation function $\bar{c}_{b}^{(2)}(z)$ and the radial distribution function $g(r)$ as shown in Fig. 2f for $\rho_{b}\sigma^{3}=0.62$. Here, we approximate $\bar{c}_{b}^{(2)}(z)$ by $(-\beta/A(\Delta z)^{2})\partial^{2}F_{\theta}/\partial\rho_{i}\partial\rho_{j}$ with $z=|z_{i}-z_{j}|$. To derive the radial distribution function, we first numerically calculate $c_{b}^{(2)}(r)$ from $\bar{c}_{b}^{(2)}(z)$ using
@@
We then obtain $g(r)$ using the Ornstein-Zernike equation. To suppress artefacts stemming from numerical transformations in the region for $g(r)\rightarrow 0$, we apply a noise-reducing filter [17].
Summarizing, our results lead to the remarkable observation that our neural free-energy functional $F_{\theta}^{(2)}$ is robust and accurate for non-uniform density systems, even far beyond its training range, solely by training on pair correlation functions in homogeneous bulk systems. These observations show that generating a large dataset of density profiles across a wide range of external potentials is not necessary to train an accurate neural free-energy functional. Instead, only a dataset of radial distribution functions from bulk systems is needed. This work thus represents a significant advancement in the field of cDFT, enabling applications to systems in arbitrary three-dimensional geometries, as accurate sampling of three-dimensional density profiles is considerably more expensive than sampling the one-dimensional radial distribution function. We expect that extensions of this method to mixtures of simple fluids and to particles with orientation degrees of freedom [14] should be possible.
Unlike the density profile, the radial distribution function is experimentally available through X-ray diffraction. This suggests the potential to use pair-correlation matching to directly construct the excess free energy of a molecular system from the true radial distribution function obtained from experiments, bypassing the dependence on simulations with model interaction approximations such as the Lennard-Jones potential.
In conclusion, this work introduces a generic machine learning approach to obtain classical free-energy functionals through pair-correlation matching. We showed how to learn a neural-network representation of the excess free-energy functional of a generic simple fluid by solely learning the direct correlations from the radial distribution function obtained from uniform density simulations. Through pair-correlation matching, we attain a neural free-energy functional that enables simultaneous and direct access to both the excess free energy and the density of a system of interacting particles in any external environment.

### v4

@@ Abstract
The intrinsic Helmholtz free-energy functional, the centerpiece of classical density functional theory (cDFT), is at best only known approximately for 3D systems. Here we introduce a method for learning a neural-network approximation of this functional by exclusively training on a dataset of radial distribution functions, circumventing the need to sample costly heterogeneous density profiles in a wide variety of external potentials. For a supercritical 3D Lennard-Jones system, we demonstrate that the learned neural free-energy functional accurately predicts planar inhomogeneous density profiles under various complex external potentials obtained from simulations.
Figure 1: 1. Bulk densities in planar geometry $\rho(z_{i})=\rho_{b}$ and radial distribution functions $g(r)$ are sampled from Monte Carlo simulations of homogeneous bulk systems of Lennard-Jones particles. Each $g(r)$ is converted to the second functional derivative of the excess free energy $\delta^{2}\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})$ by employing the Ornstein-Zernike equation. 2. Through automatic differentiation (autodiff2), the neural functional $F_{\theta}^{(2)}$ is optimized to fit the Hessian of the model output with respect to input density profiles to $\delta^{2}\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})\delta\rho(z_{j})$. 3. The optimized model can then be applied in cDFT to obtain non-uniform equilibrium density profiles through automatic differentiation (autodiff) and the free energy $F_{\theta}^{(2)}$ for a system of Lennard-Jones particles subjected to arbitrary external potentials.
@@
In this Letter, we introduce a neural free-energy functional that we train using pair-correlation matching, a novel optimization objective that matches the Hessian of the neural approximation to pair correlations of particles. We show that pair-correlation matching yields a neural functional that accurately predicts the excess free energy for a one-component system of interacting particles. The differentiable nature of this neural functional allows access to, and learning from, various structural and thermodynamic properties by utilizing the first and second functional derivatives of $\mathscr{F}_{\text{exc}}[\rho]$. Unlike previous ML approaches to cDFT [11, 12, 13, 14], our neural functional is trained by directly learning particle correlations from radial distribution functions sampled from short simulations of homogeneous bulk systems (illustrated in Fig. 1), rather than inferring them from a costly dataset of inhomogeneous densities. We demonstrate that this neural free-energy functional can be applied in the cDFT framework to achieve accurate estimates for inhomogeneous density profiles in external fields without ever having seen any inhomogeneous densities during training. Simultaneously, the neural free-energy functional provides accurate estimates of the excess free energy.
@@
with $\beta=1/k_{B}T$, $\Lambda$ the thermal wavelength, $\mu$ the chemical potential and $V_{\text{ext}}(\mathbf{r})$ the external potential. This self-consistency relation can be leveraged to find $\rho_{0}(\mathbf{r})$ through recursive Picard iteration [15, 16, 17]. Since $\rho_{0}(\mathbf{r})$ is dependent on $\mathscr{F}_{\text{exc}}[\rho]$ through the first functional derivative $\delta\mathscr{F}_{\text{exc}}[\rho]/\delta\rho(\mathbf{r})$, previous approaches to leveraging machine learning for cDFT [11, 12, 13] involve training a model to capture $\delta\mathscr{F}_{\text{exc}}/\delta\rho(\mathbf{r})$, which can be derived from sampled inhomogeneous equilibrium density profiles and employing Eq.(1).
@@
Figure 2: Evaluation of neural free-energy functionals $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$, where $F_{\theta}^{(1)}$ is optimized by matching inhomogeneous one-body densities and $F_{\theta}^{(2)}$ by pair-correlation matching in the homogeneous bulk. a) Density profiles of a Lennard-Jones system in a planar geometry characterized by an external potential (shown in gray) at a chemical potential of $\beta\mu=0$ obtained from DFT using $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ and the mean-field approximation $F_{\text{exc}}^{\text{MF}}$, along with the simulated density profile. b) Comparison of the free-energy estimates using $F_{\theta}^{(1)}$, $F_{\theta}^{(2)}$, and $F_{\text{exc}}^{\text{MF}}$ for the specific external potential shown in a), evaluated across the range $-4<\beta\mu<6$. Inset shows the bulk excess free-energy obtained from DFT and simulations across a range of bulk densities $10<\rho_{b}\sigma^{3}<$ 1. c) Mean absolute error $\frac{1}{n}\sum\left|\rho_{\text{sim}}\left(z_{i}\right)-\rho_{\text{dft}}\left(z_{i}\right)\right|$ between density profiles obtained from DFT using $F_{\theta}^{(1)}$, $F_{\theta}^{(2)}$, and $F_{\text{exc}}^{\text{MF}}$ and densities sampled from simulation. d) Absolute error of the excess free-energy from DFT and simulations. c)/d) Data is shown for 150 distinct external potentials, evaluated across the range $-4<\beta\mu<6$, with steps of $\Delta\beta\mu=0.2$. The area of the mean $\pm$ standard deviation is colored. The error is shown up to the point where the Picard iterations stop to converge to a solution within 1000 iterations. e) The bulk pressure and chemical potential obtained from DFT and simulations across a range of bulk densities $10<\rho_{b}\sigma^{3}<$ 1. f) The laterally integrated direct correlation function $\bar{c}_{b}^{(2)}(z)$ at $\rho_{b}\sigma^{3}=0.67$ and the radial distribution function $g(r)$ obtained from simulation and DFT.
@@
We train a convolutional neural network to directly learn the excess free energy $\mathscr{F}_{\text{exc}}[\rho]$, enabling the straightforward calculation of functional derivatives by (auto-)differentiating the neural functional with respect to its inputs. We focus on 3D systems in a planar geometry, where the excess free-energy of a system of area $A$ is a functional of the density $\rho(z)$, which is constant across any plane parallel to the $xy$-plane, i.e., $\rho(z)=\rho(x,y,z)$ with $\rho(x,y,z)=\rho(x^{\prime},y^{\prime},z)$ for all $(x,y),(x^{\prime},y^{\prime})$ within the confines of $A$. We represent the excess free-energy functional $\mathscr{F}_{\text{exc}}[\rho]$ as a neural network $F_{\theta}^{(2)}(\rho_{1},\cdots\rho_{n})$, with $\rho_{i}$ the density at grid point $z_{i}$ for $i\in\{1,\ldots,n\}$, network parameters $\theta$, and where the upper index ‘$(2)$’ indicates that the neural network is optimized for pair correlations of bulk systems obtained from simulations.
@@
with $\hat{h}(k)$ the Fourier transform of the total correlation function $h(r)=g(r)-1$ [8]. Thus we train our neural network such that its Hessian $\partial^{2}F_{\theta}^{(2)}/\partial\rho_{i}\partial\rho_{j}$ represents well the laterally integrated direct correlation function obtained from homogeneous bulk simulations (see End Matter for further details). We will refer to this approach as pair-correlation matching.
@@
We illustrate this methodology for systems interacting with a Lennard-Jones potential truncated at $r_{\text{cut}}=4\sigma$ with $\sigma$ the particle diameter, shifted upwards by $\epsilon_{\text{cut}}=0.98\cdot 10^{-3}\epsilon$ with $\epsilon$ the well depth, at a temperature $k_{B}T/\epsilon=2$, i.e. above the critical point. Setting $\Lambda=\sigma$ throughout, we perform grand-canonical Monte Carlo (GCMC) simulations of homogeneous bulk systems at a variety of chemical potentials $\beta\mu\in[-4,0.5]$, resulting in bulk densities $\rho_{b}\sigma^{3}\in[0.02,0.67]$. We employ $10^{9}$ trial moves in a cubic box with an edge length $L=10\sigma$ (hence with area $A=L^{2}$ and volume $V=L^{3}$) and apply periodic boundary conditions. We sample $g(r)$ and $\rho_{b}$ and convert each $g(r)$ to $\bar{c}_{b}^{(2)}(z)$ using Eqs. (2) and (3) with a grid-spacing $\Delta z=\sigma/32$. We create a data set of 800 combinations of $\bar{c}_{b}^{(2)}(z)$ and $\rho_{b}$, from which we learn the parameters $\theta$ of $F_{\theta}^{(2)}$. The neural network architecture of $F_{\theta}^{(2)}$ is a six-layer convolutional neural network with periodic padding [20] to match the system’s periodic boundary conditions. Each convolution uses a kernel size of 3 with a dilation rate of 3. The number of channels per hidden layer is configured as follows: $[16,16,32,32,64,64]$, such that $\theta$ consists of 24.4K parameters (see End Matter for further details).
@@
To evaluate the accuracy of our neural excess free-energy functional $F_{\theta}^{(2)}$, we compare it to the Van der Waals-like mean-field approximation $F_{\text{exc}}^{\text{MF}}$, which treats the attractions of Lennard-Jones particles as a perturbation on the hard-sphere system, as implemented in PyDFTlj [1]. We use the White-Bear mark II version of FMT for the excess free energy of the hard-sphere system [21]. Additionally, we compare to $F_{\theta}^{(1)}$, which is a neural free-energy functional trained by minimization of the error between $(1/\Delta z)\partial F^{(1)}_{\theta}/\partial\rho_{i}$ and $\delta\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})$ rather than by pair-correlation matching. This neural functional has the same neural network architecture as $F_{\theta}^{(2)}$ and is trained on a dataset of 800 non-uniform densities, subjected to the same set of chemical potentials as before (Supplementary Material, S. 2 [22]). Note that the free-energy functional $F_{\theta}^{(1)}$ introduced in this work is different from the neural direct correlation functional introduced by Sammüller et al. [13], since the gradient of $F_{\theta}^{(1)}$ supplies a global estimate of $\delta\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})$ for all $i\in\{1,\dots,n\}$, in contrast to a local estimate at one position $z_{i}$ (Supplementary Material, S. 8 [22]). By approximating $\delta\mathscr{F}_{\text{exc}}/\delta\rho(z_{i})$ by the gradient $(1/\Delta z)\partial F^{(n)}_{\theta}/\partial\rho_{i}$ for $n=1$ and 2, both neural functionals are applied in Picard iterations [15, 16, 17] to obtain DFT estimates for the equilibrium density profiles of inhomogeneous systems according to Eq. (1).
@@
The DFT results for an exemplary external potential at $\beta\mu=0$ are shown in Fig. 2a. We observe that the neural functionals $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ provide similar estimates that closely agree with simulation data. For the same external potential, we evaluate the accuracy of DFT estimates for the free energy for a range of chemical potentials $-4<\beta\mu<6$ (Fig. 2b). We compare with the excess free energy obtained from GCMC simulations through thermodynamic integration (Supplementary Material, S. 4 [22]). We observe that both neural functionals outperform $F_{\text{exc}}^{\text{MF}}$ within the range of $\mu$ values in the training set, exhibiting good agreement with simulations. The DFT estimates are shown until the Picard iterations diverge. We observe that $F_{\theta}^{(1)}$ diverges rapidly when extrapolating beyond the training set, even earlier than $F^{\text{exc}}_{\text{MF}}$. In contrast, $F_{\theta}^{(2)}$ converges to a solution far beyond the trained $\mu$ range.
For a more detailed comparison of the accuracy of the free-energy functionals for various inhomogeneous systems, we performed separate DFT calculations for 150 distinct external potentials, evaluated across the range $-4<\beta\mu<6$. Both $F_{\theta}^{(1)}$ and $F_{\theta}^{(2)}$ functionals show excellent agreement with simulated data for density estimates (Fig. 2c) and free-energy estimates (Fig. 2d), outperforming $F_{\text{exc}}^{\text{MF}}$ across all evaluated external potentials. Within the training range, $F_{\theta}^{(1)}$ exhibits marginally lower prediction errors than $F_{\theta}^{(2)}$, while $F_{\theta}^{(2)}$ demonstrates superior accuracy when extrapolating beyond the training range. Additional experiments show that the predictive accuracy of $F_{\theta}^{(2)}$ is more sensitive to highly inhomogeneous densities compared to $F_{\theta}^{(1)}$, though this effect is reduced when incorporating higher bulk densities into the training set of $F_{\theta}^{(2)}$ (Supplementary Material, S. 7 [22]). Although we might expect that training on bulk data provides significantly less information about the underlying functional than training on inhomogeneous data, these results show that pair-correlation matching works remarkably well in practice.
@@
and the chemical potential $\mu$ following Eq. (1). Again, we find excellent agreement of both neural functionals with simulations within the training set, and superior agreement of $F_{\theta}^{(2)}$ at higher densities, as shown in Fig. 2e. Lastly, we demonstrate that $F_{\theta}^{(2)}$ compared to both $F_{\theta}^{(1)}$ and $F^{\text{exc}}_{\text{MF}}$ provides accurate estimates for the laterally integrated direct correlation function $\bar{c}_{b}^{(2)}(z)$ and the radial distribution function $g(r)$ as shown in Fig. 2f for $\rho_{b}\sigma^{3}=0.67$. Here, we approximate $\bar{c}_{b}^{(2)}(z)$ by $(-\beta/A(\Delta z)^{2})\partial^{2}F_{\theta}/\partial\rho_{i}\partial\rho_{j}$ with $z=|z_{i}-z_{j}|$. To derive the radial distribution function, we first numerically calculate $c_{b}^{(2)}(r)$ from $\bar{c}_{b}^{(2)}(z)$ using
@@
We then obtain $g(r)$ using the Ornstein-Zernike equation (Eq. 3). To suppress artefacts stemming from numerical transformations in the region for $g(r)\rightarrow 0$, we apply a noise-reducing filter (Supplementary Material, S. 5 [22]).
Our results lead to the surprising observation that the neural free-energy functional $F_{\theta}^{(2)}$ is robust and predicts accurate non-uniform density profiles, even substantially beyond its training range and solely by training on pair correlation functions of homogeneous bulk systems. This training on homogeneous bulk systems offers an alternative to existing training schemes of classical DFT, which so far have all been based on training on density profiles of inhomogeneous fluids in a variety of external fields [11, 12, 13, 14]. For pairwise systems, this alternative could be particularly useful if pair-correlations are the only data available (e.g. from scattering experiments on homogeneous fluids), if the external potential leads to computationally expensive 3D density profiles, or if spatial heterogeneity is coupled to orientational anisotropy (e.g. in liquid crystals, non-spherical molecules, or patchy particles [14]). The presently proposed training on homogeneous bulk systems can also be a computationally inexpensive addition to the existing training protocols on inhomogeneous fluids, where an optimal combination may be found between training on homogeneous and inhomogeneous states. In this sense, the present study can be seen as an extreme case of training on pair correlations in the bulk only.
In conclusion, this work introduces a generic machine learning approach to obtain classical free-energy functionals through pair-correlation matching, through which we attain a neural free-energy functional that enables simultaneous and direct access to both the excess free energy and the density of a classical supercritical system of interacting particles in any external environment.